### v4

# The physical basis of teleological evolutions

Abstract
It is easily seen that the mere existence of an optimal quantum computational speedup logically implies the mutually exclusive $or$ of well-defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. Involving only the measurements of commuting observables, quantum algorithms are submitted to the same (classical) logic and should therefore host the quantum superposition of the causal loops in question. However, being causal in character, their ordinary quantum description cannot describe causal loops. We argue that it must therefore be incomplete and show that it is completed by time-symmetrizing it. This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by the quantum superposition of causal loops in question. In their completed quantum description, quantum algorithms respond to the pre-scientific notion of teleological evolution, that is, an evolution toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again). Once applied to the evolutions of the living, this notion was dismissed with the advent of modern science for the alleged absence of a physical basis. We show that, under a quantum cosmological interpretation of the Anthropic Principle, the same quantum superposition of causal loops underlying the teleological character of quantum algorithms becomes the missing physical basis of the teleological character of natural evolutions.
1 Introduction
The notion of teleological evolution is that of an evolution toward a goal with an attractor in the same goal it will produce in the future. It is, of course, a puzzling notion since it necessarily implies a causal loop. As is well known, it was often applied to the evolutions of the living before the scientific revolution. With the advent of the latter, it was dismissed because of its alleged lack of physical basis. That was, of course, under the notion of causal evolution that is dominant in science even today.
We take the opposite position in the present work. We will show that, in quantum mechanics, there is a well defined physical basis to the notion of teleological evolution.
In the first place, we show that quantum algorithms are teleological evolutions, indeed evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again).
Of course, seeing this requires a quantum description of quantum algorithms different from the ordinary one, which is naturally that of a causal process. The latter quantum description, with causality always going in the same time-direction, cannot describe causal loops.
We show that the ordinary quantum description of quantum algorithms, although complete from the mathematical standpoint, is physically incomplete and is completed by time-symmetrizing it. This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by a quantum superposition of causal loops. In each of them, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps). Of course, we are dealing with an evolution that responds to the notion of teleological evolution: it is indeed toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again).
This completion of the quantum description provides a unified quantitative explanation of all the optimal quantum speedups achieved in solving oracle problems. Given an oracle problem, the number of computation steps needed to solve it in an optimal quantum way is that needed to solve it in a optimal logical way benefiting of the advanced knowledge of any of the possible halves of the information about the solution. Of course, there is no such an explanation in current quantum computing science. Obviously, we have a causal quantum description for each quantum algorithm, but there is no known way of unifying them.
This part of the work is a further development of the retrocausal approach to the quantum computational speedup $\left[1-5\right]$.
Of course, both the approach in question and the part of the work discussed so far are unconventional and potentially controversial. The notion of causal loop is puzzling and mostly extraneous to today’s physics. However, we believe that this time we have given a very simple and easily checkable demonstration that the quantum computational speedup is physically due to quantum causal loops.
Let us outline it. We consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem. If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[6\right]$ – she always solves the problem by opening just one drawer222 In the four drawers case, Grover algorithm yields the solution with certainty. In the following, by Grover algorithm we will always mean Long’s $\left[7\right]$ version of it, which does that always. . This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of viewing it.
It should already be clear from the above that that of the quantum speedup is a fundamental discovery that changes a long standing way of viewing things333 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[8]$ in 1969. In 1982, Feynman $\left[9\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources. In 1985, Deutsch $\left[10\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup. .
We can go now to our demonstration, which is surprisingly simple. Of course, we must adopt a way of thinking open to retrocausality.
The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers, in a mutually exclusive way one of the three pairs obtained by pairing the drawer with the ball with another drawer. In fact, if the ball is in the drawer opened, she has solved the problem. If not, it must be in the other drawer and she has solved the problem of locating the ball as well, having opened just one drawer.
Then the question becomes: from where does Alice always get the information that the ball is in a given pair of drawers? Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. This is obviously a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution by opening correspondingly fewer drawers. This, in the mutually exclusive $or$ of all the possible ways of taking half of the information about the solution.
We would argue that, with the above elementary logic, we have, so to speak, broken the sound barrier: the quantum speedup does logically imply causal loops. Then, one gos from logic to physics as follows.
Involving only the measurements of commuting observables (the problem-setting and the corresponding solution), quantum algorithms are submitted to the same (classical) logic. Therefore they must host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of an optimal quantum speedup; of course, logical mutual exclusivity of possibilities, moving to the quantum world, becomes their quantum superposition. We have already seen that the quantum description of this quantum superposition is generated by time-symmetrizing the ordinary quantum description of the quantum algorithm. This time-symmetrization must therefore be performed because, as we will see, is the only way of generating the causal loops that must be hosted in the quantum algorithm.
Let us add a few comments to the above.
By ordinary quantum description of quantum algorithms, we do not mean their usual quantum description. The latter is limited to the unitary process that solves the problem, where the quantum speedup resides, and the final measurement of the solution. It is trivially incomplete since it lacks the initial measurement. Since we need to complete the quantum description, in the first place we should emend this trivial incompleteness. We do this by extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings.
In the case of Grover algorithm, we obtain a unitary evolution between an initial and a final one-to-one correlated measurement outcomes encoding respectively the problem-setting (the number of the drawer with the ball selected by the initial measurement) and the corresponding solution of the problem. We have called it the ordinary quantum description of the quantum algorithm, ordinary since it is still a causal description that cannot describe causal loops. Note that any quantum algorithm that produces the solution with certainty can be put in this form. If the solution is not an invertible function of the problem-setting, it suffices that Alice, at the end, measures both the solution and the problem setting. This is a trivial but important step: of course, the usual quantum description of quantum algorithms would completely obscure the approach we are following.
Let us eventually note that we are dealing with the same kind of incompleteness of the ordinary quantum description that was ascribed by Einstein, Podolsky and Rosen $\left[11\right]$ to the quantum description of quantum nonlocality. Also this quantum description can be seen as a unitary evolution (the process of spatial separations of the two subsystems) between two one-to-one correlated measurement outcomes (those of measuring in the same basis the two respective maximally entangled observables) – we must think that this process goes from one to the other measurement outcome via the time the two subsystems were not spatially separate. Time symmetrized, it generates a quantum superposition of causal loops that play the role of the hidden variables envisaged by those authors to physically (locally) explain quantum non-locality $\left[5\right]$.
This ends the first part of the work, a review and further clarification of the retrocausal approach to the quantum speedup.
Of course, under the above explanation of their quantum speedup, quantum algorithms turn out to be teleological evolutions. The second part of the work is about the possibility that the quantum causal loops that stand at the basis of the teleological character of quantum algorithms are the missing physical basis of the teleological character of the evolutions of the living.
This immediately raises two problems.
One is explaining the presence of quantum features (the quantum speedup) in evolutions that can be the size of the Earth and have durations of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels points to the only possible solution: to scale them down from the largest and longest existing evolution there is, that of the universe. Of course, at the same time, we should adopt quantum cosmology, that is, the assumption that the universe – a perfectly isolated system since by definition there is nothing outside of it – evolves in a unitary quantum way.
The other is to find a formal analogy between quantum algorithms and, under quantum cosmology, natural evolutions. Obviously, it should justify the fact that also natural evolutions can occur with a quantum speedup and are consequently teleological in character.
Interestingly, the Anthropic Principle comes to our rescue. The Fine-tuned Universe version of it asserts that the unitary evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today, in particular (“sentient life” or “no life”). The reason for this correlation is that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
Having adopted quantum cosmology, of course, also implies a quantum interpretation of the Anthropic Principle. Simply, the correlation between the value of the physical constants since the beginning of time and the value of life today becomes quantum and its building up is performed by the universe’s unitary quantum evolution.
This is what we were looking for. In fact, it is exactly what quantum algorithms do. They unitarily build up a quantum correlation between the problem-setting since the beginning of them and the solution of the problem at their end. Under this formal analogy, the evolution of the universe toward the value of life could be endowed with quantum speedup and would even in the case that the quantum search for the correlation in question is random, exactly as in random quantum search algorithms $\left[12\right]$.
Of course, being endowed with quantum speedup, under the quantum causal loops explanation of it, the evolution in question is teleological in character. In the universe we are in, what becomes teleological in character is the evolution of the universe toward sentient life.
Eventually, as we shall see, the evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out everything else, even if decohered, should inherit the quantum speed and thus the teleological character of the latter evolution. The same would consequently hold for any evolution of life on Earth, e. g. that of species.
The notion of a teleological evolution of life is of course puzzling. For example, the Darwinian theory of evolution would change as as follow: those random DNA mutations that, in the future, will survive natural selection would become more probable. With the past influenced by the future, this is a completely new way of viewing the physical world. No wonder the teleological notion has been rejected by science.
In conclusion, the work is highly unconventional, but its findings were obtained in a surprisingly simple and well-verifiable way. If it were correct, as we think, it would be a paradigm shift not only for the natural sciences, but also for the way we see physical reality.
2 Rationale
As the present work is about a fundamentally different way of viewing the role of causality in quantum (correlation) processes, before going to the formalizations we provide a Rationale that describes it conceptually.
This Rationale is divided into two parts corresponding to the two parts in which the work is divided.
The first provides the concept of the time-symmetric quantum causal loops, showing how they stand at the physical basis of the quantum computational speedup and quantum nonlocality. In particular, the time-symmetric quantum causal loop description of quantum algorithms highlights their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution of the problem again).
The second addresses the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
2.1 The time-symmetric quantum causal loops – the concept
We go now to the first part of the work. We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[6\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. Of course, we must ignore the ban on the use of the notion of retrocausality.
The fact that Alice (the problem-solver) always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
Indeed, in any of these three mutually exclusive cases, she opens either drawer; if the ball is in it, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well.
The question is then: from where does Alice always get (in a mutually exclusive way) this advanced knowledge of one of the possible halves of the information that specifies the solution? Since, to her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. This is, of course, a causal loop that allows her to produce the solution by opening just one drawer. There must be, in a mutually exclusive way, one causal loop for each of the three possible ways of knowing in advance one of the possible halves of the information that specifies the solution.
We consider now $N$ drawers. Grover algorithm (also in Long’s version) yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case. Also now, this quantum speedup is logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is in one of the possible tuples of $\sqrt{N}$ drawers. She then solves the problem, in the best possible logical way, by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
We can therefore formulate the following advanced knowledge rule:
The number of computation steps (drawer openings) needed to solve Grover’s problem in an optimal quantum way is that needed to solve it in an optimal logical way given the advanced knowledge of one of the possible halves of the information about the solution.
Of course, involving only the measurements of commuting observables (the problem-setting and the corresponding solution), Grover algorithm, as any other quantum algorithm, is submitted to the same (classical) logic. Its quantum description should therefore describe the quantum superposition of the causal loops in question (logical mutual exclusivity of possibilities becomes their quantum superposition in the quantum world).
However, the above immediately raises a problem. The current quantum description of quantum algorithms, as that of any quantum process, is causal in character. In other words, causality always goes in the same time direction. Obviously, it is a description that cannot describe causal loops; we tentatively ascribe this to its incompleteness.
Before going to this assumed incompleteness, of a fundamental kind, we should emend a trivial incompleteness of the current quantum description of quantum algorithms. It is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is obviously incomplete because it lacks the initial measurement. Then, we extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings. In the case of Grover algorithm, where the solution is an invertible function of the problem-setting (both are the number of the drawer with the ball), we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement). Of course, this unitary evolution never changes the problem-setting selected by the initial measurement.
By the way, seeing quantum algorithms as unitary evolutions between an initial and a final one-to-one correlated measurement outcomes is a key step of the present work. In view of going beyond Grover algorithm, we note that, if the solution is not an invertible function of the problem-setting, we can have one-to-one correlation between the initial and final measurement outcomes as well by requiring that, at the end, the problem-solver measures both the solution and the problem setting. The only constraint is that the quantum algorithm determines the solution with certainty.
We call the above, less incomplete, quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops. We should complete this ordinary quantum description in a way that leaves it mathematically unaltered but changes its causality structure – describes the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup. This is simply done time-symmetrizing it as follows.
Instead of ascribing the selection of the pair of one-to-one correlated measurement outcomes all to the initial measurement, as customary, we require that it time-symmetrically shares between the initial and final measurements, of course in a quantum superposition of all the possible ways. In each time-symmetrization instance, the initial measurement selects one of the possible halves of the information that specifies the sorted out pair of correlated measurement outcomes (it selects a quantum superposition of problem-settings instead of the customary single one). Its measurement outcome unitarily propagates forward in time until becoming the state immediately before the final measurement, a quantum superposition of tensor products each a problem-setting multiplying the corresponding solution. The final measurement completes the selection, of course by selecting a single pair problem-setting and corresponding solution. Its measurement outcome, obviously the same of the ordinary quantum description, time-symmetrically propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous initial measurement outcome.
This backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm. Note that, mathematically, i.e. disregarding the time-direction of causality, it is the ordinary quantum description back again.
Also note that, together, the forward and backward in time propagations describe a causal loop that changes its own past (we call it time-symmetric since it is generated by a time-symmetrization procedure). As seen by the problem-solver, from whom the outcome of the initial measurement should always be concealed (would give her information about the solution – see Subsection 4.2.), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops. Of course, the backward in time propagations of all time-symmetrization instances are identical with each other. Therefore, as required, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered. However, it changes the structure of causality along it. As seen by the problem-solver, the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; then, she can use this knowledge to produce the solution with fewer computation steps. If she exploits it entirely, the quantum speedup is optimal.
In equivalent terms, in each causal loop, the computational complexity of the problem to be solved reduces – this will become tangible in the formalizations of Section 4. The number of computation steps needed to solve the reduced complexity problem in an optimal quantum way is that needed to solve it in an optimal logical way. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces.
In view of what will follow, it is important to note that these results have been obtained independently of the distinctive character of the unitary evolution between the two correlated measurement outcomes (i.e. the actual quantum algorithm). It sufficed to know that it is a unitary evolution that one-to-one correlates two initially unrelated observables (the problem-setting observable and the solution observable) never changing the eigenvalue of the observable selected by the initial measurement.
In the present work, we will always know that there is such a unitary evolution. It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[4\right]$. Under the assumption that between two one-to-one correlated measurement outcomes there can always be a unitary evolution of the kind in question, the advanced knowledge rule holds for any oracle problem.
Eventually, we note that the time-symmetrization in question, and the consequent time-symmetric quantum causal loops, are imposed by logic and quantum mechanics. Summarizing, logic says that the sheer existence of an optimal quantum speedup logically implies the mutually exclusive $or$ of well defined causal loops. Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup that, involving only the measurements of commuting observables, is submitted to the same (classical) logic. In turn, it is easily seen that the only way of obtaining these causal loops is by time-symmetrizing the ordinary quantum description of the quantum algorithm. In fact, to have a causal loop, we must properly share the selection of the pair problem-setting and corresponding solution between the initial and final measurements. Because of the time-symmetry of the unitary evolution between the two one-to-one correlated measurement outcomes, we should share it evenly. This exactly specifies the time-symmetrization we are dealing with. Therefore, as long as logic and quantum mechanics are consistent, the time-symmetric quantum causal loops generated by the time-symmetrization operation should be exempt from the paradoxes that plague the usual kind.
Interestingly, the time-symmetrization we are dealing with turns out to be mandatory also to satisfy a symmetry implicit in any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement.
In fact, in this kind of process, we can postpone the initial measurement at will – all results must remain unaltered. When the two measurements become simultaneous, the time-symmetrization in question must already be there to become the atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
This further motivation for time-symmetrizing the unitary evolutions we are dealing with, besides further justifying the previous motivation, makes things much more general. Any such unitary evolution must be time-symmetrized, what leaves it mathematically unaltered but changes its ordinary, causal, quantum description into that a quantum superposition of time-symmetric causal loops.
If the unitary evolution is a quantum algorithm that determines the solution with certainty, in each causal loop it is as if the problem solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future. However, this just creates the room for its occurring with quantum speedup; the quantum algorithm does not need any more to be optimal like in the case of Grover algorithm. How much of this room is exploited depends on the problem-solver, i.e. on the quantum algorithm.
If the unitary evolution in question is not a quantum algorithm, but anyhow creates one-to-one quantum correlation between two initially unrelated observables (without changing the eigenvalue of the observable selected by the initial measurement), it can always occur with quantum speedup. In this more general case, the problem-solver advanced knowledge might loose any meaning but, of course, can be replaced by the reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place. This reduction, in turn, creates the room for a quantum speedup.
There is also the case that the unitary evolution does not build up any quantum correlation. It is the case of quantum nonlocality, namely the measurement in the same basis of two maximally entangled observables whose respective subsystems are in a state of spatial separation. Here maximal entanglement between the two observables is already there before the space-separation process. After the time-symmetrization, this process becomes a quantum superposition of time-symmetric quantum causal loops locally going from one to the other measurement outcome via the time the two subsystems were not yet spatially separate. Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[5\right]$. We can see that these causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e. also physically) describe quantum nonlocality.
Note that, as a consequence, the incompleteness we are currently ascribing to the quantum description of quantum algorithms is the same that Einstein and the others ascribed to the quantum description of quantum nonlocality. In both cases, what is incomplete is the quantum description of a quantum correlation. With this, we end the first part of the work.
2.2 Teleological character of natural evolutions – a possible physical basis
The second part of the work examines the possibility that the time-symmetric quantum causal loops underlying the teleological character of quantum algorithms are also the missing physical basis of the teleological character of the natural evolutions for which the teleological notion was conceived.
This, of course, immediately raises the problem of explaining the existence of quantum features – indeed the quantum speedup – in evolutions that can be the size of the Earth and have a duration of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Naturally, at the same time we should adopt quantum cosmology $\left[13,14\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[15-19\right]$, the “Fine-tuned Universe” version of it. This version of the Principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental physical constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”. Let us assume, for simplicity, that this correlation today is one-to-one. The significance of this principle lies in the fact that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
Under a quantum cosmological interpretation of it, the correlation addressed by the Anthropic Principle becomes quantum. We should think of a value of the fundamental constants observable and a value of life observable. At the beginning, there should be a quantum superposition (or mixture) of all the possible universes, each with its value of the fundamental constants. An initial measurement of the value of the fundamental constants observable would select one value at random. The subsequent unitary evolution of the universe would produce, at the present time, the corresponding value of life, which would then be measured. In Subsection 5.2., after the formalization of the notion of the time-symmetric quantum causal loops, we will provide a formalized toy model of the building up of the correlation in question on the part of the universe’s evolution.
By the way, since the value of the fundamental constants never changes along the universe’s evolution, in the universe we are in we can do without the initial measurement (when there were no observers). Today, there would be a quantum superposition of tensor products, each a possible value of the fundamental constants multiplying the corresponding value of life. The measurement of either observable today would select the eigenvalues of both maximally entangled observables. Everything is as if the value of the fundamental constants was selected at the beginning of time.
Summing up, we are dealing with a unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue selected by the initial measurement (here the value of the fundamental constants). We have seen that this unitary evolution can always occur with quantum speedup – there is the room for it. Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[12\right]$. It should consequently be teleological in character.
In the universe we are in, the search for one-to-one correlation between the value of the fundamental constants and the value of life becomes search for sentient life. It is thus the universe’s evolution toward sentient life that occurs with quantum speedup and is consequently teleological in character.
In principle, the evolution of life on Earth can be obtained from that of the universe toward sentient life by tracing out all the rest. Evolution times would be identical and the amount of the resources employed by the evolution of life on Earth could not exceed those employed by its image in the evolution of the universe that comprises them. In the end, the evolution of life on Earth, even if decohered, would occur with quantum speedup and would thus be teleological in character. This should also hold for the evolutions of species, etc..
Of course, all the above is puzzling.
A physical basis to the dismissed notion of teleological evolution would be a paradigm shift for the natural sciences. As is well known, empirically, the teleological notion better fits with the functional structure and the evolutions of the living. If it has been discarded, it is only because of the assumed absence of a physical basis. For the time being, the fact that quantum algorithms are certainly teleological evolutions, and that, under a quantum cosmological interpretation of the Anthropic Principle, natural evolutions would be teleological in character too, may shake the current belief that such a physical basis cannot exist.
That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the usual strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence. Things become quite different in the case of the evolutions of the living. The fact that what happens around and within us would be driven not only by the past but also by a future even billions of years ahead is more than puzzling. It might also go along with ancient beliefs, but of a kind deliberately ignored by science. Once science was accused of heresy for denying a physical basis to beliefs of the kind in question. We could be facing the reverse problem now: proposing a physical basis to the notion of teleological evolution could well be seen as a scientific heresy.
3 Positioning the work
We try and position the present work. Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Besides the retrocausal approach to the quantum computational speedup $\left[1-5\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
The present approach is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of being open to the violation of the ordinary notion of causality, not in common the way of looking at the quantum description of quantum algorithms. For example, in $\left[27\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time. So far so good but, consequently, they are excluded (by the approach in question) from the phenomena that can accommodate causality violations. Of course, for us, this means putting too much faith in their current quantum description. As we have seen, with their quantum computational speedups, quantum algorithms should be a most clear example of violation of one-way causality – what obviously implies changing their current quantum description.
Some concepts of the work have precedents.
One is the local explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate. Here, still locally, it would tell the other observable the state to assume in view of its future measurement. This retrocausal, local, explanation of quantum nonlocality is similar to ours. The difference is that sending backward in time an entire measurement outcome can originate paradoxes and, of course, is not imposed by quantum mechanics.
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[30\right]$ for an account of this. This tells us a lot on the physicist’s position toward retrocausality, even at the top levels. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[13\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the value of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating from before Deutsch’s 1985 discovery, the two principles in question make no reference to the quantum speedup of the universe’s evolution and, a-fortiori, its consequent teleological character.
4 The time-symmetric quantum causal loops – formalization
We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[6\right]$ quantum search algorithm. As anticipated in the Rationale, this immediately generalizes first to any number of drawers, then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. The formalization consists of the following steps:
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem. We call this extended quantum description the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
(2) This point is propaedeutic to point (4). It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary external observer. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her information about the solution of the problem.
(3) Complete the ordinary quantum description of the quantum algorithm by time-symmetrizing it. Mathematically, the quantum description of the process between the two correlated measurement outcomes remains unaltered. However, the description of the behavior of causality along it changes: the single causal process of the ordinary quantum description is replaced by a quantum superposition of time-symmetric causal loops.
(4) Relativize the quantum description of the causal loops generated at point (3) with respect to the problem-solver. This produces the quantum description of the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
4.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
The usual quantum description of quantum algorithms is limited to the process of solving the problem, where the quantum speedup resides, and lacks the initial measurement. Of course, it is trivially incomplete: the canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement. We extend the usual quantum description to the upstream process of setting the problem. An initial measurement of the problem-setting, performed in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random. This is followed by the unitary evolution that computes the corresponding solution and the final measurement of the solution.
Let us introduce the notation. We make reference to the simplest four drawer instance of Grover algorithm:
We number the four draws in binary notation: $00,01,10,11$.
We will need two quantum registers of two quantum bits each:
A register $B$ under the control of the problem-setter Bob contains the problem-setting (the number of the drawer in which the ball is hidden). The content of register $B$ is the problem-setting observable $\hat{B}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{B}$ and $00$; $|01\rangle_{B}$ and $01$; $|10\rangle_{B}$ and $10$; $|11\rangle_{B}$ and $11$.
A register $A$ under the control of the problem-solver Alice contains the number of the drawer opened by her. The content of register $A$ is the observable $\hat{A}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$;
we call it content of $A$ observable. Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
The quantum description of the extended quantum algorithm, namely its ordinary quantum description is:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table I)
Here and in the following we disregard normalization.
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow444 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operantion that changes nothing for us. .
It is now Alice’s turn. The unitary part of her problem-solving action, $\hat{\mathbf{U}}_{1,2}$, produces the solution of the problem, of course, never changing the problem-setting. Note that, of $\hat{\mathbf{U}}_{1,2}$, it is sufficient to know that it does this, whereas we will never need to know how it does it. Of course, we must also know that there exists a unitary transformation that does this, Grover algorithm in the present case.
At the end of Alice’s problem-solving action, at time $t_{2}$, register $A$ contains the solution of the problem, i.e. the number of the drawer with the ball – bottom-right corner of the diagram.
Eventually, by measuring $\hat{A}$ (the content of $A$), which is already in one of its eigenstates, Alice acquires the solution with probability $1$ – the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement. Correspondingly, there is a unitary transformation between the initial and final measurement outcomes; the process between them is reversible despite comprising the final measurement.
4.2 Relativizing the ordinary quantum description of the quantum algorithm with respect to the problem-solver
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process. It is not a description for the problem-solver (Alice). In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement555 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution. Knowledge of the problem-setting selected by the initial measurement would tell her also the solution. .
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by Bob and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following description of the quantum algorithm with respect to Alice:
time $\begin{array}[]{ccc}\begin{array}[]{c}\text{time}\text{ }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{\text{time }\text{}}t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ &&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table II)
To Alice, the initial measurement of $\hat{B}$ leaves the initial state of register $B$ unaltered. At the beginning of her problem-solving action she is completely ignorant of the number of the drawer with the ball selected by Bob – see the state of register $B$ on the left of $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$.
The unitary transformation $\hat{\mathbf{U}}_{1,2}$ is then performed for a quantum superposition of the four possible numbers of the drawer with the ball. The state at the end of it – on the right of the horizontal arrows – is a quantum superposition of tensor products, each a possible number of the drawer with the ball in register $B$ tensor product the corresponding solution (that same number) in register $A$. Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution – bottom-right corner of Table II diagram.
Note that this relativization relies on relational quantum mechanics $\left[33,34\right]$, where the quantum state is relative to the observer. In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
4.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
As we have already said, in order to obtain the quantum superposition of the causal loops whose mutually exclusive $or$ is implied by the sheer existence of an optimal quantum speedup, we should first time-symmetrize the ordinary quantum description of the quantum algorithm, then relativize with respect to the problem-solver the causal loops obtained.
To start with, we should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes among all the possible pairs time-symmetrically shares between the initial and final measurements. The half information selected by the initial measurement should propagate forward in time, by $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement. Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time, by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$, until replacing the previous outcome of the initial measurement.
We consider the time-symmetrization instance where the initial measurement of $\hat{B}$ reduces to the measurement of $\hat{B}_{l}$, the left bit of the two-bit number contained in register $B$; say that it randomly selects left bit 00. The final measurement of $\hat{A}$ should correspondingly reduce to that of $\hat{A_{r}}$ the right bit of the two-bit number contained in register $A$; say that it randomly selects right bit $1$. The randomly selected number of the drawer with the ball is thus $01$.
In this particular time-symmetrization instance, the ordinary quantum description of the quantum algorithm with respect to the external observer of Table I becomes:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, }\text{}\text{meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }\text{}t_{2}\text{, meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}\\ &&\Downarrow\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table III)
The initial measurement of $\hat{B}_{l}$, selecting the 00 of $01$, projects the initial quantum superposition of all the basis vectors of register $B$ on the superposition of those beginning with 00 – vertical arrow on the left of the diagram. Under $\hat{\mathbf{U}}_{1,2}$, the latter superposition causally propagates forward in time into the superposition of the two tensor products on the right of the right looking horizontal arrows. Then the final measurement of $\hat{A_{r}}$, selecting the $1$ of $01$, projects the latter superposition on the term ending with $1$ under the right vertical arrow. Note that this outcome of the final measurement, which inherits both selections, coincides with that of the ordinary quantum description – see the bottom-right corner of Table I diagram. Time-symmetrically, this final measurement outcome causally propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation (which inherits both selections) is an instance of the time-symmetrized description of the quantum algorithm – bottom line of the diagram. Note that, since the symbol “$\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$” is mathematically equivalent to “$\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$”, this instance, up to the time-direction of causality, is the unitary process between the two one-to-one correlated measurement outcomes back again.
Also note that causality performs a causal loop that changes its own past; the outcome of its initial measurement of $\hat{B}_{l}$ is changed into the outcome of the initial measurement of $\hat{B}$ of the ordinary quantum description; compare the state of register $B$ under the left vertical arrow with the state in the bottom-left corner of the diagram.
We should eventually take the quantum superposition of all the time-symmetrization instances. We can see that, no matter how we evenly share the selection between initial and final measurements, the final measurement outcome, which inherits both selections, is always that of the final measurement of $\hat{A}$ of the ordinary quantum description of the quantum algorithm. Correspondingly, its subsequent backward in time propagation by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ (bottom line of the diagram), mathematically (disregarding the direction of causality), is always the ordinary quantum description of the unitary process between the two correlated measurement outcomes, namely the bottom line of Table I diagram.
Summing up, mathematically, time-symmetrization leaves the unitary process between the two measurement outcomes unaltered but changes the structure of causality along it. The causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
4.4 Relativizing the causal loops with respect to the problem-solver
The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is relative to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup. This is because the latter causal loops, concerning the problem-solver advanced knowledge of one of the halves of the information about the solution, must of course be relative to the problem-solver.
Therefore, we should relativize with respect to the problem-solver the causal loop of Table III by postponing after her problem-solving action the projection of the quantum state due to the initial measurement of $\hat{B}_{l}$ (whose outcome should be concealed from Alice). This yields the causal loop in the following table:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }t_{2}\text{,meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ \begin{array}[]{c}\left(|01\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$ – top-left corner of the diagram – is postponed after Alice’s problem-solving action, namely outside the present table that is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in present assumptions selects the value $1$ of the right bit of the number contained in register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized description of the quantum algorithm with respect to the problem-solver.
Note that we would have obtained the identical diagram by time-symmetrizing the description of the quantum algorithm relativized to the problem-solver of Table II.
Also note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, the bottom line of the diagram of Table IV can also be read in the usual left to right way:
time $\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table V)
Note that the entire causal loop is the zigzag diagram of table IV. It is a causal loop that changes its own past. The initial state in the top-left part of the diagram, where the state of register $B$ tells that Alice is completely ignorant of the number of the drawer with the ball, changes into the initial state in the bottom-left part of the diagram, where the state of register $B$ tells that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$. Since what happens along the time-symmetric causal loop obeys classical logic, the number of drawer openings needed to find the solution in the optimal quantum case is that needed in the optimal logical way, namely just one.
In other words, the computational complexity of the problem to be solved by her reduces. The problem of locating the ball in one of $4$ drawers – as for the state of register $B$ in the upper-left corner of the diagram of Table IV – becomes that of locating it in one of $2=\sqrt{4}$ drawers – as for the state of register $B$ in the bottom-left corner of Table IV diagram (or the left of Table V diagram). In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $4$ basis vectors to $2=\sqrt{4}$.
In equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces.
The generalization of the above is straightforward. Note that, in all our considerations, we never needed to know the unitary transformation between the two measurement outcomes $\hat{\mathbf{U}}_{1,2}$. It sufficed to know that it unitarily changes its input state, where the two observables ($\hat{B}$ and $\hat{A}$ ) are unrelated, into its output state, where they are one-to-one correlated, without changing the eigenvalue of the observable selected by the initial measurement.
First, the causal loops explanation of the quantum speedup, and the fact that the latter is quadratic in kind, remain there if we increase the number of quantum bits in the two registers of an even number and in an equal way. In fact, all our considerations remain identical. This means, in particular, Grover algorithm with number of drawers any even power of $2$.
The same, of course, holds, for any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. This will apply to the teleological evolutions addressed in the next section.
5 The physical basis of the notion of teleological evolution
We have seen, by the way rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in the optimal logical way benefiting of the advanced knowledge in question. In equivalent terms, the computational complexity of the problem to be solved or the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces. As already observed, in quantum search algorithms, also a random quantum search for the solution is endowed with (a suboptimal) quantum speedup.
Therefore, we can say that quantum algorithms are evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future. Of course, this is necessarily a causal loop and responds to the notion of teleological evolution. As is well known, before the scientific revolution, the evolutions of the living were often considered to be teleological in character. The teleological notion was then dismissed with the advent of modern science for the alleged absence of a physical basis. That was, of course, under the causal view of evolutions that is dominant in science also today.
Now, the fact that quantum algorithms turn out to be teleological evolutions should already cast a doubt on this way of seeing. It also, naturally, raises the question whether the time-symmetric quantum causal loops that stand at the basis of their teleological character could be the missing physical basis of the teleological character of natural evolutions.
As already seen in the Rationale, under a quantum cosmological interpretation of the Anthropic Principle, the search for the value of life by the evolution of the universe under the fundamental physical laws and constants has the same form a quantum search algorithm. This search can therefore occur with quantum speedup and does even if it is random. Consequently, the evolution of a universe toward the value of life turns out to be teleological in character. Of course, in the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
Eventually, as we have seen, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within the quantum cosmological view.
This section is a discussion of the many impacts that the present teleological view of natural evolutions might have on different areas of science and knowledge; of course, not forgetting the natural philosophy that gave rise to the notion of teleological evolution.
5.1 Decohered evolutions of life endowed with quantum speedup: how would they look?
All natural evolutions are necessarily described in a probabilistic way. We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical stochastic process, transition probabilities can depend on the past of the process but not on the future of it.
Instead, a decohered process endowed with quantum speedup must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection must make, in the present, the transition toward them more probable. By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the process would be difficult to distinguish from a classical stochastic process. Since also in this case the time-symmetric causal loops we are dealing with should be involved, let us call time-symmetric also the just outlined kind of stochastic evolution.
We mention what could be an emblematic example of time-symmetric stochastic evolution. It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
The usual explanation is of course a causal one: at first, feathers appeared through random mutations and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
According to the present teleological view of the evolutions of the living, even during the flightless period of the feathered dinosaurs, those random DNA mutations that would have led in the future to the conquest of the air – obviously an extremely successful outcome from the point of view of survival – would have more likely occurred.
The above is puzzling, of course. We should keep in our mind that we are under the quantum cosmological assumption that might make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as strange as the microscopic quantum world.
5.2 Teleological evolution of the universe toward life – a toy model
A toy model of a teleological evolution of the universe toward life can be trivially mapped on the simplest instance of Grover algorithm.
We should imagine a value of the fundamental physical constants observable $\hat{B}$ and a value of life observable $\hat{A}$. Say that the eigenvalue $01$ of $\hat{B}$ is the value of the fundamental physical constants compatible with the development of sentient life and the eigenvalues $00,10,11$ the values of those incompatible. We assume that, correspondingly, the eigenvalue $01$ of the measurement of $\hat{A}$ at the present time would mean existence of sentient life and the eigenvalues $00,10,11$ absence of life.
Initially, $\hat{B}$ should be in a quantum superposition (or, indifferently, mixture) of all its eigenstates (i.e of all the possible universes each with its fundamental physical constants). We can assume that also $\hat{A}$, the value of life observable, is initially in a quantum superposition of all its eigenstates.
The universe’s unitary evolution under the fundamental physical laws and constants would have finally, at the present time, built up a quantum superposition of tensor products, each an eigenstate of $\hat{B}$ tensor product the corresponding eigenstate of $\hat{A}$ – see table II. In our universe, at the present time one with observers, the measurement of $\hat{A}$ at the present time would select the eigenvalue $01$ of $\hat{A}$, namely “existence of sentient life” and the corresponding eigenvalue $01$ of the value of the fundamental physical constants observable $\hat{B}$. Since the universe’s evolution does not change the value of the fundamental physical constants, all is as if the eigenvalue of $\hat{B}$ was selected by a measurement of $\hat{B}$ at the beginning of time, when there could be no observers.
Of course, this is exactly the quantum search algorithm described in the previous section, the version with respect to the problem-solver of Table II. As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup, and does even if the quantum search for the value of sentient life is random.
5.3 How to model a teleological natural evolution
We address the problem of how to model a decohered evolution endowed with quantum speedup – say that of the living on Earth after the beginning of the liquid water stage of it. The canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest. Of course, in the present case, this holds only in principle. However, it would be interesting to apply the canonical way to a toy model of the situation we are dealing with.
One should identify the purely quantum evolution of a system (the universe) endowed with quantum speedup, where the achievement of the goal is mostly localized in a subsystem thereof (the Earth). Then, one should obtain a decohered evolution of the subsystem from that of the system by tracing out all the rest. Note that the interaction of the subsystem with the rest of the system, in the present case, would do the opposite than decoherence. It would be what allows the evolution of the subsystem toward the goal to occur with quantum speedup as part of the purely quantum evolution of the system.
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious (approximate) way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[35\right]$.
As already said, a further possibility should be to model a decohered evolution endowed with quantum speedup as a stochastic process with transition probabilities time-symmetrically depending on both the past and the future of the process, something just outlined for the time being.
5.4 Impact on the natural sciences
The current position of the natural sciences toward teleology can be considered in some measure conflicting. On the one hand, the teleological view would seem to allow a deeper understanding of the functional structure and the evolutions of living beings. On the other, as this vision is considered without physical basis and therefore misleading, there is outright hostility to it.
This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[36\right]$:
Teleology is like a mistress to a biologist: he cannot live without her, but he is unwilling to be seen with her in public.
It would seem that, today, no evolutionary biologist supports the teleological view, some explicitly assert that it is an unworkable view because it implies retrocausality.
No biologist, of course, does not mean no scientist. A paradigmatic example is the American philosopher Nagel $\left[37\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
We would put together Nagel’s view in support of the teleological character of the evolutions of the living and Haldane’s assertion that a biologist cannot live without teleology to affirm that the notion of teleological natural evolution is more “scientific” than the mainstream one of causal natural evolution. This in the sense that it better answers empirical evidence – of course the first canon of science. The only reason biologists would be against teleology in spite of empirical evidence would come from another science. We mean the notion – current in today’s physics – that a teleological evolution is physically impossible.
Of course, the sheer notion that quantum computation is teleological in character may already put a strain on the position of today’s biology: actually, although in the quantum world, teleological evolutions are physically possible. The strain should increase with the notion that, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward life should be teleological in character. It should further increase with the consequent notion that also the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, even if decohered, should inherit the teleological character of the universe’s evolution.
Summing up, given the above, the impact of the present teleological view of natural evolutions on today’s evolutionary biology would seem to be almost a head on collision.
Almost because, on the one side, in recent years a group of evolutionary biologists $\left[38\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a way out.
On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[35\right]$.
5.5 Impact on philosophy
Of course, the current theory of teleological evolutions would rehabilitate an ancient notion of natural philosophy $\left[39\right]$. We should therefore be allowed to look for the possibility that there are other relations between this theory and philosophy. As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
In effect, the teleological theory in question seems to revamp a full body of ancient philosophical views sort of opposite (and complementary) to the body of philosophical views emerged with the scientific revolution.
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
The causal view and the mutual causality view seem to have respectively subtended two opposing views throughout the history of philosophy. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views respectively.
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[13\right]$, and the present theory is a mutual causality view.
We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond to the causal and mutual causality views of evolutions, respectively.
For what concerns reductionism vs monism, we note that deriving the teleological character of the decohered evolutions of the living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view. It presupposes that what we see around and within us is essentially inseparable from the whole.
Let us mention other theories in the history of philosophy reminiscent of the mutual causality between sentient life today and the physical laws with their physical constants since the beginning of time:
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, are the causes of natural phenomena and constitute their unity” $\left[40\right]$. In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. The identity between the fundamental physical laws since the beginning of time and the ideas of our mind today is formally reminiscent of the identity (and mutual causality) between the number of the drawer with the ball (the problem-setting) randomly selected by Bob’s measurement at the beginning of the quantum algorithm and the number of the drawer with the ball (the solution of the problem) selected by Alice’s measurement at the end of it. It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, when it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
The above should also go along with Fritjov Capra’s $\left[41\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
(ii) We do not need to be anthropocentric. This mutual causality would also be between physical laws and spatial intelligence, often superior to our in non-human animals. The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality). Indeed a quantum speedup in the development of spatial intelligence would imply the mutual causality in question.
(iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[42\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
(iv) It is natural to wander whether the present time-symmetric quantum causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
(v) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[43\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today? Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[44\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it? We would not feel like dismissing these analogies as sheer coincidences. As a matter of fact, in idealistic philosophy – in Schelling – aesthetic insight would reproduce the fundamental mechanisms of the production of life.
6 Research prospects
Future work on the subject of the present work could have a broad spectrum.
For what concerns quantum physics:
(i) Look for a synthesis between the present notion of the time-symmetric quantum causal loops and time-symmetric quantum mechanics $\left[20-24\right]$ to which it is obviously inspired.
(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (the mostly autonomous evolution of a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
(iii) The mechanism described in (ii) might apply to other life processes for which the presence of a quantum behavior is suspected, for example, to the possible quantum behavior of the brain. Although mostly autonomous from the rest of the universe, this behavior could derive from that of a quantum universe by tracing out everything external to the brain. Note the radical difference between this view and the usual one where people looks for an autonomous quantum behavior in the macroscopic, warm and wet brain. By the way, we could introduce a distinction between two kinds of quantum behaviors: autonomous, like those of microscopic systems sufficiently isolated from all the rest, and derived, namely quantum behaviors of even macroscopic subsystems obtained from that of a quantum universe – the system – by tracing out all the rest.
(iv) Look for the possible impact of the teleological view on quantum cosmology. According to current assumptions, even the inanimate part of the latter, always an evolution toward life, should be endowed with quantum speedup and thus have teleological character.
(v) The Anthropic Principle is limited to stating the correlation between the “value” of life and the value of the fundamental physical constants since the beginning of time. One could also look for a correlation between the value of life and the character of the fundamental physical laws since the beginning of time. Of course, also these would have an attractor in the goal of creating life.
About the natural sciences, looking for:
(I) A reconsideration of the ban on applying the teleological view and language to natural evolutions. Restoring the dismissed notion of teleological evolution would obviously be a paradigm shift for today’s natural sciences.
(II) Toy models of teleological Darwinian processes. These evolutions could be either approximated as pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
(III) An application of the time-symmetric quantum causal loops to quantum self-organization. An example is studying the teleological formation of a self-replicating system. There seem to be an enormous jump of complexity in the transition from a soup of life-precursor molecules into a self-replicating system. The idea that this transition has an attractor in the full fledged self-replicating system can therefore be attractive.
Finally, opposing schools of thought throughout the history of philosophy could be reviewed in the light of the opposition between the causal and mutual causality views, as exemplified in the previous subsection.
7 Conclusions
The work is, of course, highly unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
We have (trivially) shown that the sheer existence of the optimal quantum speedup of Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic. Therefore, Grover algorithm must host the quantum superposition of the causal loops in question.
However, being causal in character as that of any other quantum process, its customary quantum description cannot describe causal loops; we ascribe this to its incompleteness.
Before going to the way of remedying this incompleteness, we must eliminate another, trivial, one. The customary quantum description of quantum algorithms is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is incomplete since it lacks the initial measurement.
Thus, as a first step, we extend the customary quantum description of the process of solving the problem to the upstream process of setting the problem, by adding the initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings. We have thus a unitary evolution between two one-to-one correlated measurement outcomes that does not change the problem-setting selected by the initial measurement. This is a key feature of quantum algorithms, not visible in their customary quantum description777 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting. . We call this kind of process the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
The second step is to change the causal structure of the ordinary quantum description of the quantum algorithm by time-symmetrizing the latter. We time-symmetrically share between the initial and final measurements the selection of the information that specifies the sorted out pair of correlated measurement outcomes, in a quantum superposition of all the possible ways.
Each time-symmetrization instance, mathematically, yields the unitary evolution between the two measurement outcomes back again, but the structure of causality along it changes. It becomes a causal loop that changes its own past. Relativized with respect to the problem-solver (who should be concealed from any outcome of the initial measurement), the causal loop in question becomes one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops. Of course, the time-symmetrization leaves the ordinary quantum description mathematically unaltered but changes its single causal process into a quantum superposition of causal loops.
Until now, the time-symmetrization in question has been imposed by the necessity of reproducing, in the quantum description of Grover algorithm, the causal loops logically implied by the sheer existence of its (optimal) quantum speedup. However, as we have seen, the same time-symmetrization is also redundantly imposed by a symmetry implicit in any unitary evolution between an initial and a final, one-to-one correlated, measurement outcomes that does not change the eigenvalue selected by the initial measurement. As a consequence, any such evolution is a quantum superposition of time-symmetric causal loops.
If the unitary evolution builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement, like in the case of Grover algorithm, in each time-symmetrization instance the dimension of the Hilbert space in which the search for one-to-one correlation takes place reduces. This creates the room for a quantum speedup. There is a quantum speedup even in the case that the search for one-to-one correlation is random.
The process is a quantum superposition of time-symmetric causal loops also in the case that the two observables are already maximally entangled at the beginning of it. It is the case of quantum nonlocality, where the quantum process starts with the spatial separation between two subsystems whose respective observables are already maximally entangled. Here the causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell either observable the state to assume in view of its future measurement. They play the role of the hidden variables envisaged by Einstein and the others to locally explain quantum nonlocality.
The quantum causal loop description of quantum algorithms highlights their teleological character, namely their being evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again). This naturally raises the question whether the time-symmetric quantum causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
This immediately raises the problem of explaining a quantum speedup in evolutions of the living that can be the size of the Earth and have durations of billions of years. The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the only possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time we should adopt quantum cosmology, namely the assumption that the universe evolves in a unitary quantum way.
We should now call into play the “Fine-tuned Universe” version of the Anthropic Principle. It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”. We assume for simplicity that this correlation is one-to-one.
With the evolution of the universe quantum, also the correlation asserted by the Anthropic Principle becomes quantum. We are thus dealing with a unitary evolution that one-to-one correlates two initially unrelated observables (the value of the fundamental constants and the value of life observables) without changing the eigenvalue of the observable selected by the initial measurement. We have seen that this kind of unitary evolution can occur with quantum speedup and does so even in the case that the search for one-to-one correlation is random.
The evolution of the universe toward the value of life should therefore take place with quantum speedup and thus be teleological in character. In the universe we are in, this would apply to the evolution of the universe toward sentient life.
We have also seen that the evolution of life on Earth, in principle obtainable from that of the universe toward sentient life by tracing out all the rest, should inherit the teleological character of the latter evolution. In fact: (i) evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe and (ii) the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
In conclusion, the notion of the time-symmetric quantum causal loops turns out to be the common physical basis of such diverse physical phenomena as the quantum computational speedup, the consequent teleological character of quantum algorithms, quantum nonlocality, and – under a quantum cosmological interpretation of the Anthropic Principle – the teleological character of the evolutions of the living.
Of course, that of teleological evolution is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
Regarding the quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
Regarding the teleological character of the macroscopic evolutions of life, the situation becomes even more puzzling. We limit ourselves to noticing that this is a necessary consequence of the quantum cosmological assumption. Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience some of the strange characters of the microscopic quantum world.
Acknowledgments
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments throughout the whole of the present approach to the quantum speedup, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the Elsag Bailey-ISI Turin Villa Gualino workshops of the years 1991-96.