### v7

# The physical basis of teleological evolutions

Abstract
It is easily seen that the mere existence of an optimal quantum computational speedup logically implies the mutually exclusive $or$ of well-defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. Involving only the measurements of commuting observables, quantum algorithms are submitted to the same (classical) logic and should therefore host the quantum superposition of the causal loops in question. However, being causal in character, their ordinary quantum description cannot describe causal loops. We argue that it must therefore be incomplete and show that it is completed by time-symmetrizing it. This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by the quantum superposition of causal loops in question. In their completed quantum description, quantum algorithms respond to the pre-scientific notion of teleological evolution, that is, an evolution toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again). Once applied to the evolutions of the living, this notion was dismissed with the advent of modern science for the alleged absence of a physical basis. We show that, under a quantum cosmological interpretation of the Anthropic Principle, the same quantum superposition of causal loops underlying the teleological character of quantum algorithms becomes the missing physical basis of the teleological character of natural evolutions.
1 Extended summary
In this work we propose a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the same goal it will produce in the future. It is, of course, a puzzling notion since it implies a causal loop. For centuries it has been applied by natural philosophy to the evolutions of the living; it was then dismissed with the advent of modern science because of its alleged lack of a physical basis.
What we think should change that (also centuries old) position of science is the relatively recent discovery of the quantum computational speedup. In this work, by quantum (computational) speedup we mean its strict form, that is, the fact that quantum algorithms solve problems with fewer logical operations (computation steps) than classically possible – sometimes, the quantum speedup is with respect to the best known classical algorithm but one does not whether it is over what is classically possible.
As we try to be interdisciplinary, let us provide an example of (strict) quantum speedup right away. Consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem. If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of seeing the thing.
Let us mention priorities along the path that led to this fundamental discovery. The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[2]$ in 1969. In 1982, Feynman $\left[3\right]$ highlighted the higher computational efficiency of the quantum world, showing that the classical simulation of a quantum process can require an amount of physical resources exponentially higher than the amount involved in the quantum process itself. In 1985, Deutsch $\left[4\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
The reason we believe that that position of science should be changed is simply the fact that quantum algorithms turn out to be teleological evolutions, that is, evolutions toward a goal (the solution of the problem) with an attractor in the same goal they will produce in the future (the solution again). We will show that this is the immediate consequence of a retrocausal approach to the quantum speedup of ours and co-authors $\left[5-9\right]$.
We chose the extended summary form because of the unconventionality of the work: there are many concepts that need to be explained from the beginning. The work is divided into two parts. The first is a review and further clarification of that retrocausal approach to the quantum speedup: it will justify the teleological character of quantum algorithms. The second is whether the physical basis of the teleological character of quantum algorithms could also be the physical basis of a teleological character of the evolutions of the living for which the teleological notion was conceived. The present introduction is divided into two corresponding parts.
1.1 The teleological character of quantum algorithms
Let us focus on Grover quantum search algorithm, by the way a most universal quantum algorithm that can solve with a quadratic quantum speedup any NP problem.
We show that just opening to retroausality allows to produce a simple quantitative explanation of the optimal quantum speedup of Grover algorithm that will also apply to a very large class of quantum speedups.
Let us begin with the four drawers case. Our argument goes as follows.
The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before she begins her problem-solving action, that the ball is in a given pair of drawers and benefiting of that knowledge to solve the problem by opening just one drawer. In fact, if the ball is in the drawer opened, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. Of course, what Alice should know in advance is the mutually exclusive $or$ of the three possible ways of pairing the number of the drawer with the ball with that of another drawer. Note that this also means knowing in advance half of the information that specifies the solution of the problem.
Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. Of course, this implies retrocausality, more specifically a causal loop.
The above immediately generalizes to any number, $N$, of drawers. Grover algorithm yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case. Also now, this quantum speedup can be logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is (in a mutually exclusive way) in one of the possible tuples of $\sqrt{N}$ drawers. She then solves the problem in the best possible logical way by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
We can conclude that the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of well defined causal loops, each for each possible way of taking half of the information about the solution. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution by opening fewer drawers.
We can formulate the following advanced knowledge rule that, as checked in $\left[8\right]$, also holds for all the major quantum algorithms for the solution of oracle problems:
The number of computation steps (oracle queries) needed to solve the (oracle) problem in an optimal quantum way is the number needed to solve it in an optimal logical way benefiting of the advanced knowledge of one (any one) of the possible halves of the information about the solution.
Later we will show why this rule applies to any problem for which a problem-setting and a corresponding solution can be identified. By the way, it is a rule that does not exist in mainstream quantum computing science. There, the only way of knowing the number of computation steps needed to solve a problem with a quantum speedup is finding a quantum algorithm that does that. As is well known, this is a matter of high ingenuity. Of course, this goes along with the fact that mainstream quantum computing science, as mainstream quantum mechanics, relies on the notion of causal evolution (i.e., with causality always going in the same time-direction) that is incompatible with that of causal loop.
Let us make a digression. As the work goes on, we will naturally go into more and more detail, with the consequent risk of masking the forest with the trees. So, we would like to emphasize that the forest is right now in front of our eyes. The notion of causal loops is obviously controversial; probably most people believe that causal loops cannot exist except perhaps in rather exotic physical situations. Now we would have, moreover trivially, demonstrated that the quantum speedup of a quantum algorithm logically implies their existence. This is the forest! Since this will have puzzling consequences, we suggest keeping in mind this fact that the very existence of the quantum speedup implies causal loops. Of course, it is the causal loops that are the source of all the weirdness in question.
Back to our line of thinking, we should ask ourselves what is the impact of the fact that a quantum speed up logically implies the mutually exclusive $or$ of those causal loops on our way of seeing the quantum algorithm, that is, on the quantum description that we make of it.
To this end, we should make reference to a quantum description of the quantum algorithm less incomplete than the usual one, which is limited to the process of solving the problem (where the quantum speedup resides) and lacks the initial measurement. Of course, the canonical quantum description of a quantum process must include the initial and final measurements and the unitary evolution in between.
Naturally, this problem is easily adjusted. It suffices to extend the usual quantum description of the process of solving the problem to the upstream process of setting the problem. The extended quantum description is thus as follows. An initial measurement, in a quantum superposition (or mixture) of all the possible problem settings, selects one of them at random. This is followed by the unitary computation of the corresponding solution and by the final measurement that reads the solution without changing the quantum state222 With four drawers, Grover algorithm yields the solution of the problem with certainty; when the drawers are more than four, by Grover algorithm we will always mean Long’s version of it $\left[10\right]$, which always provides the solution with certainty. . Note that there is one-to-one correlation between the initial measurement outcome, which encodes the problem-setting, and the final one, which encodes the corresponding solution – in Grover algorithm, they are both the number of the drawer with the ball. Furthermore, there is a unitary evolution connecting the two measurement outcomes, despite the fact it comprises the final measurement (which indeed does not change the quantum state). This extension of the usual quantum description of quantum algorithms to the process of setting the problem will play a key role in the present work. Let us call it the ordinary quantum description of the quantum algorithm, ordinary because it is still the ordinary, causal, quantum description of mainstream quantum mechanics.
We can go now to the impact of the logical implication in question on our way of describing the quantum algorithm.
The first thing to observe is that, involving only the measurements of commuting observables (the problem-setting and the solution observables), the quantum algorithm is subject to the same (classical) logic of the logical implication in question; in other words, the non-classical quantum logic of non-commuting observables is not involved here. The quantum algorithm that delivers the optimal quantum speedup must therefore accommodate the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its quantum speedup (in the quantum world, the logical mutual exclusive $or$ of possibilities becomes their quantum superposition). However, being causal in character, its ordinary quantum description cannot describe causal loops. According to the present line of thinking, we must conclude that the ordinary quantum description of the quantum algorithm is either wrong or incomplete, since it does not and, being causal, cannot describe causal loops. We will see that it is incomplete as it is completed by time-symmetrizing it.
We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes – the problem-setting and the corresponding solution. The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement, that performed by the final measurement, time-symmetrically, should propagate backward in time until replacing the previous initial measurement outcome. This is a causal loop that changes its own past; it is, indeed, one of the causal loops logically implied by the sheer existence of the quantum speedup. Of course, we should then take the quantum superposition of all the possible ways of evenly sharing. We will see that all this, mathematically, leaves the description of the unitary part of the quantum algorithm unaltered but changes that of the behavior of causality along that part. The single, unitary, causal process of the ordinary quantum description of the quantum algorithm changes into the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
We will (easily) show that this time-symmetrization is the only possible way to generate the quantum superposition of the causal loops in question, which in turn must be hosted within the quantum algorithm. Therefore, according to the present line of thinking, it is imposed by logic and quantum mechanics.
Most importantly to present ends, the fact that, in each causal loop, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution with fewer computation steps highlights the teleological character of quantum algorithms. They are indeed evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (the solution again).
Until now we have focused on Grover quantum search algorithm, but the results obtained go well beyond it.
In the first place, we never needed to know the unitary evolution between the two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (i.e., Grover algorithm). It sufficed to know that such an evolution exists.
In the second place, until now, the time-symmetrization of the ordinary quantum description of the quantum algorithm has been imposed by the necessity of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its quantum speedup. However, it turns out that there is another reason for time-symmetrizing, independent of this.
Let us consider the unitary (thus reversible) part of the quantum algorithm. It is a unitary evolution interconnecting two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement. This allows to postpone the initial measurement at will. When made simultaneous with the final measurement, the time-symmetrization in question must already be in place to become the perfect atemporal symmetry between the simultaneous measurements of two maximally entangled observables.
The two things together make the results obtained in the case of Grover algorithm applicable to a very general class of quantum processes involving quantum correlation. In fact, they imply that:
The complete quantum description of any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing its ordinary quantum description.
We call the above the causal loop rule of quantum correlation.
We can see that it is a generalization also when the process is still a quantum algorithm. In fact, until now, all the results obtained are for an optimal quantum algorithm, Grover algorithm. However, also the time-symmetrization of the ordinary quantum description of non-optimal quantum algorithms yields a quantum superposition of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. The point is that this time she is free to exploit any amount of this information. For example, this covers the case of the random quantum search algorithm $\left[11\right]$, which provides a suboptimal quantum speedup. We will be interested in this kind of quantum algorithms further on.
In view of applying the causal loop rule of quantum correlation to processes that have the form of quantum algorithms but are not that, let us note that we can do without the anthropomorphic notion of a problem-solver that produces the solution with fewer computation steps because she knows in advance half of the information about it; we can also avoid to speak of “quantum algorithm”. We can say that, in each time-symmetrization instance, there is a reduction of the dimension of the Hilbert space in which the search for the one-to-one correlation between the initial and final measurement outcomes takes place and consequently the correlation in question can be achieved with a quantum speedup.
Of course, the process does not even need to have the form of a quantum algorithm. In other words, it does not need to create one-to-one correlation between two initially unrelated observables. The two observables can be maximally entangled before the process begins. As we shall see, quantum nonlocality is exactly a process of this kind and the quantum superposition of causal loops generated by its time-symmetrization explains the EPR (Einstein, Podolsky, and Rosen) paradox – it provides a local explanation of quantum nonlocality.
1.2 The possible teleological character of the evolutions of the living
The teleological character of quantum algorithms naturally raises the question of whether the quantum superposition of causal loops that underlie it can also be the missing physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived. Having established that, although in the quantum world, there is a physical basis for the notion of teleological evolution, probing this possibility could even be considered a duty of the physical science toward the natural sciences.
One difficulty, of course, would be to explain the presence of quantum features – the quantum speedup – in the evolutions of life that can be Earth-sized and last billions of years. However, the obvious impossibility that the quantum features of the microscopic world scale up to such levels points to another possibility: that they scale down from the largest and longest existing evolution, that of the universe. Of course, at the same time, we should adopt quantum cosmology $\left[12,13\right]$, i.e., the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way.
Surprisingly, the corresponding quantum cosmological interpretation of the Anthropic Principle $[14-18]$ brings all things together. The Fine-tuned Universe version of this principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today (in particular: “presence of sentient life” or “absence of life”). As is well known, this principle stems from the fact that the slightest variation in the value of the fundamental constants would have produced a radically different universe, most likely incapable of hosting life. Let us assume, for simplicity, that this correlation is one-to-one.
The quantum interpretation of the Anthropic Principle produces a formal identity between the building up, by the unitary evolution of the universe under the fundamental physical laws and constants, of the quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today on the one side and, on the other, the building up, by a quantum search algorithm, of the quantum correlation between the value of the problem-setting since its beginning and the value of the solution at its end. In other words, the unitary evolution of the universe toward the value of life, formally, turns out to be a quantum search algorithm. These quantum algorithms yield a quantum speedup even if the search for the correlation is random $\left[11\right]$. It is indeed reasonable (and conservative) to assume that the search for this correlation by the universe’s evolution under the fundamental physical laws and constants – its search for life in the universe we are in – is random.
According to this formal identity, the evolution of the universe toward the value of life should occur with quantum speedup and would consequently be teleological in character. In the universe we are in, the same should hold for the evolution of the universe toward life. It is then easy to show that this teleological character should be inherited by the evolution (and the evolutions) of life on Earth, derivable from that of the universe by tracing out everything else. In fact, evolution times should be the same for the evolution of life on Earth and its image in the evolution of the universe and the resources employed by the former evolution cannot exceed those involved by the latter that comprises them.
Summing up, coherently with the present line of thinking, we should conclude with the conditional statement that, if the universe evolved in a unitary quantum way, the evolutions of the living would be teleological in character.
In the next Section 2, we try to position the present work within the literature that challenges the still dominant notion of causal evolution. In Section 3, we review and further develop the quantum causal loop description of quantum algorithms. Section 4 is the (of course, speculative) transposition of the teleological description of quantum algorithms to natural evolutions. Section 5 is on the prospects for further research. Section 6 gives the conclusions.
2 Positioning the work
Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[19-23\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[22\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[24-28\right]$.
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and, of course, has in common with the attempts in question the fact of looking beyond the ordinary notion of a causality that always goes in the same time-direction, not in common the approach. In fact, these attempts explicitly rule out the possibility that quantum algorithms, with their causally ordered quantum circuits, can accommodate causality violations $\left[26,27\right]$. For us instead, with their quantum speedup over what is classically possible, they are the paradigmatic example of causality violation. Evidently, the two approaches address different kinds of causality violations.
Some concepts of the work have precedents.
One is the celebrated EPR paper $\left[29\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description is incomplete. More specifically, they claimed that the quantum description of quantum nonlocality is incomplete because it mathematically predicts something that it does not physically explain, namely the “spooky action at a distance”. As we will see, time-symmetrizing the ordinary quantum description of quantum nonlocality, as we did with that of quantum algorithms, completes it. It generates a quantum superposition of causal loops that play the role of the hidden variables those authors envisaged to complete the quantum description.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[30\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
This retrocausal explanation of quantum nonlocality is similar to ours as it also involves a causal loop. The difference is in the character of the causal loops. Sending backward in time an entire measurement outcome can originate paradoxes and, of course, is not imposed by quantum mechanics but is instrumental to finding a local explanation of quantum nonlocality. As already said, the causal loops originated by time-symmetrizing the ordinary quantum description are instead imposed by quantum mechanics. Therefore, first, they must be there independently of the fact of explaining quantum non-locality locally; second, as long as quantum mechanics is consistent, they should be exempt from the paradoxes that plague the usual kind.
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[31\right]$ for an account of this. This tells us a lot on the physicist’s position toward retrocausality, even at the top levels. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[32,33\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating to before Deutsch’s 1985 discovery, of course the two principles in question make no reference to a quantum speedup of the universe’s evolution and its consequent teleological character.
3 Completing the quantum description of quantum algorithms
This section is divided into two parts. The first begins with a constructive critique of the ordinary quantum description of Grover algorithm and ends with a generalization of the results obtained. We show (i) why the quantum description in question, being causal in character, must be considered incomplete, (ii) how it should be completed, and (iii) how both things extend to a very general class of quantum correlation processes. The second part formalizes the completion procedure consisting in a time-symmetrization of the ordinary quantum description.
3.1 Updating the EPR critique of the quantum description
Of course, the first assertion that the quantum description is incomplete goes back to the famous EPR paper by Einstein, Podolsky and Rosen $\left[29\right]$. As is well known, Einstein in particular, argued that the ordinary quantum description of quantum nonlocality is incomplete because it does not physically explain the “spooky action at a distance” that it mathematically predicts. They also stressed the need to devise ”hidden variables” that would complete the quantum description.
The present critique of the quantum description could be considered an update of that of Einstein and the others because:
(i) it shows that Einstein et al.’s assertion that the quantum description of quantum nonlocality is incomplete can be extended to quantum algorithms, of course discovered exactly half a century after the publication of the EPR paper;
(ii) the reason for the incompleteness of the quantum description turns out to be the same in both cases;
(iii) in the quantum computational context, it turns out to be evident that the quantum description should be completed by time-symmetrizing it: completing in the same way the quantum description of quantum nonlocality, locally (physically) describes the spooky action at a distance.
First, we develop our critique of the quantum description on that of the four drawers instance of Grover quantum search algorithm $\left[1\right]$. For convenience, we summarize the part of this critique anticipated in the introduction.
In its simple instance, the problem solved by Grover algorithm is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. In the quantum case – by the optimal quantum search algorithm of Grover – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case.
It is natural to ask ourselves how it is possible that, in the quantum world, a ball hidden in one of four drawers can always be located by opening a single drawer. Of course, if one goes into the details of the quantum algorithm, one sees the ingenious interplays between quantum state superpositions and interferences that lead to producing the location of the ball by opening just one drawer. However, these interplays are naturally different from quantum algorithm to quantum algorithm and, as far as we know, in mainstream quantum information there is no known way of unifying a quantitative reason for which they yield a quantum speedup. A unified, quantitative, reason is provided in the retrocausal approach to the quantum speedup of ours and coauthors but, as things are now, this unconventional approach never took off within the scientific community.
In the next subsection we address the unification problem. We will see that a unifying quantitative explanation of the quantum speedups becomes quite evident once the quantum descriptions of quantum algorithms is complete. It is a unifying explanation since it is the same for all quantum algorithm. Since deriving it requires an unconventional way of reasoning, open of course to retrocausality, we will give first the rationale then, in subsection 3.2., the formalizations.
3.1.1 Grover quantum search algorithm as a quantum superposition of causal loops
We have already shown in the Introduction that the sheer existence of the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps). In the optimal quantum case, the number of computations steps is that needed to solve the problem in an optimal logical way benefiting of the advanced knowledge in question.
Always in the introduction, we wandered what was the impact of the above fact on the current quantum description of the quantum algorithm, noticing that one should take a quantum description not limited to the process of solving the problem but extended to the upstream process of setting the problem.
Our ordinary quantum description of the quantum algorithm therefore became:
An initial measurement in a quantum superposition (or mixture) of all the possible problem-settings selects one of them at random. This is followed by the unitary computation of the corresponding solution, which of course does not change the problem-setting selected by the initial measurement, and the final measurement of the solution, which does not change the state immediately before it.
We called this quantum description ordinary because it is still the customary, causal, quantum description that cannot describe causal loops.
We observed that, involving only the measurements of commuting observables (the problem-setting and the solution), quantum algorithms are submitted to the same (classical) logic of the logical implication in question. Therefore, Grover algorithm should host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its optimal quantum speedup. Of course, this is not represented by its present ordinary quantum description, but we show that it is after completing this quantum description by time-symmetrizing it.
This time-symmetrization is, of course, submitted to the requirement of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. As we shall see, the only possible way of obtaining this quantum superposition is by requiring that the initial and final measurements time-symmetrically contribute to the selection of the sorted out pair of one-to-one correlated measurement outcomes.
Things should go as follows. In each time symmetrization instance, the initial measurement selects one of the halves of the information that specifies the sorted out problem-setting (and thus also the corresponding solution) – instead of selecting a single problem-setting as customary, it selects a suitable quantum superposition of problem-settings. Its measurement outcome unitarily propagates forward in time becoming, immediately before the final measurement, a quantum superposition of tensor products, each one of the problem-settings selected by the initial measurement multiplying the corresponding solution. The final measurement of the solution completes the selection. Note that, encoding the tensor product of the sorted out problem-setting and the corresponding solution, the final measurement outcome is the same of the ordinary quantum description. Eventually, time-symmetrically, this measurement outcome propagates backward in time, by the inverse of the time-forward unitary evolution, until replacing the previous initial measurement outcome.
At this point, we should make two important observations:
(i) The backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the quantum algorithm. Disregarding the time-direction of causality, read from left to right, it is the unitary part of the ordinary quantum description of the quantum algorithm back again. Since this is the same in each time-symmetrization instance, also the quantum superposition of all instances is that unitary part.
Summing up, we can say that the time-symmetrization leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered. What was incomplete is the physical part of the quantum description, that is, the description of the behavior of causality along the unitary part of the quantum process.
(ii) In each time-symmetrization instance, the forward and the backward in time propagations, together, constitute a causal loop that changes its own past (the outcome of the initial measurement). As we will see in the formalizations of subsection 3.2, this also changes the initial complete ignorance of the solution on the part of the problem-solver into knowledge of one of the possible halves of the information about it. There is a corresponding reduction of the computational complexity of the problem to be solved and the dimension of the Hilbert space in which the search for the solution takes place. The number of computation steps needed to solve the problem in an optimal quantum way is the number needed to solve in an optimal logical way the reduced computational complexity problem. We are of course dealing with one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
It is easy to show that the above time-symmetrization is the only possible way of generating the causal loops whose quantum superposition the quantum algorithm must host.
In fact, to have a causal loop, in the first place the selection of the information that specifies the sorted out pair of correlated measurement outcomes must properly share between the initial and final measurements with the selection performed by the initial measurement propagating forward in time until becoming the state immediately before the final measurement and that successively performed by the final measurement propagating backward in time until replacing the previous initial measurement outcome. Of course, this creates the forward and backward in time parts of a causal loop. Moreover, to (exactly) generate the causal loops logically implied by the sheer existence of the optimal quantum speedup, the sharing should be even.
Note that this is nothing but Laplace’s demon brought into the quantum world. Here, due to the possibility of having quantum state superpositions, there is more freedom. Indeed, it is a freedom exploited by the time-symmetrization process to create the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
Eventually, let us observe that the quantum causal loops we are dealing with are generated by a mandatory time-symmetrization procedure – mandatory to generate the causal loops that must be hosted in the quantum algorithm. Therefore, as long as quantum mechanics is consistent, these causal loops should be exempt from the paradoxes that plague the causal loops of the usual kind. When needed to distinguish them from the latter kind, we will call them time-symmetric causal loops because generated by a time-symmetrization procedure.
3.1.2 Generalization
All the above has been developed for Grover algorithm. However, let us show how it generalizes. We should make two premises:
(i) All the results achieved until now have been obtained with no need of knowing the unitary evolution between the two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement – i.e. the quantum algorithm itself. We only needed to know that such a unitary evolution exists. This was the case until now, the unitary evolution being that of Grover algorithm (it will always be the case in the following).
(ii) Since the eigenvalue selected by the initial measurement remains unchanged throughout the unitary evolution in question, one is free to postpone that measurement at will. When performed simultaneously with the final measurement, the time-symmetrization we are dealing with must already have been put in place to represent the perfect (atemporal) symmetry between the simultaneous measurements of two maximally entangled, one-to-one correlated, observables. This is a second motivation for time-symmetrizing the ordinary quantum description of the quantum process in question, independent of the need of reproducing in the quantum algorithm the causal loops logically implied by the sheer existence of the quantum speedup. There is, of course, a redundancy between the two motivations. We produced both because together they say more.
Putting (i) and (ii) together, brings us well beyond the optimal Grover algorithm considered until now. In fact we obtain that:
The complete quantum description of any unitary evolution between an initial and a final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing it.
We call the above the causal loop rule of quantum correlation. Let us examine a few of its consequences.
In the first place, to be a quantum superposition of causal loop, the quantum algorithm does no more need to be optimal333 If the solution is not an invertible function of the problem setting, we can have one-to-one correlation between the initial and final measurement outcomes by requiring that the final measurement measures both the solution and the problem-setting. . The time-symmetrization of its ordinary quantum description always yields that it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. In equivalent terms, there is a corresponding reduction of the computational complexity of the problem to be solved, and the dimension of the Hilbert space in which the search for the solution (or, identically, the one-to-one correlation) takes place. However, any of these facts just creates the room for a quantum speedup. To also have a quantum speedup, it takes that the problem-solver, or more generally the unitary evolution, exploits such a room. The room is fully exploited in the optimal quantum case, but it can also be not exploited at all or only partly exploited.
The latter is the case of the random quantum search algorithm $\left[11\right]$ that yields a suboptimal quantum speedup. Later, we will need to make reference to this quantum algorithm.
Of course, the quantum process may not even be a quantum algorithm. For example, quantum nonlocality is a process of this kind.
It in fact consists of a unitary evolution (the process of spatial separation between the two subsystems) between two one-to-one correlated measurement outcomes – those of measuring in the same basis the two maximally entangled observables. The process of spatial separation, of course, never changes the eigenvalue selected by any measurement. Here, each causal loop locally goes from one to the other measurement outcome via the time the two subsystems were not yet spatially separate where, always locally, it changes the state of both observables in a way that ensures the appropriate correlation between the two future measurement outcomes – see $\left[9\right]$. This is exactly the role that Einstein and the others ascribed to the hidden variables they envisaged.
Thus, the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and the others, in their (EPR) paper, ascribed to the quantum description of quantum nonlocality. Of course, in both cases what is incomplete is the ordinary quantum description of a quantum correlation.
By the way, as is well known, the EPR paper inspired works of great fundamental and practical interest like Bell’s inequality $\left[34\right]$ and Ekert’s quantum cryptography based on Bell’s theorem $\left[35\right]$. We do not understand why the argument concerning the incompleteness of the quantum description, central in that paper, has been removed by the scientific community. As is also well known, the question of the incompleteness of the quantum description has been at the center of the famous Einstein-Bohr dispute. Bohr, who was for the completeness, was eventually decreed the winner by the scientific community. Apparently that was once and for all.
According to the present line of thinking, also this position of science should be changed. To us, of course, Einstein should have been the winner. Another time, he would have put his finger on a very sensitive point in the quantum mechanical theory.
Eventually, as we shall see in Section 4., under a quantum cosmological interpretation of the Anthropic Principle, also the evolution of the universe toward the value of life turns out to be a unitary evolution between two one-to-one correlated measurement outcomes (the value of the fundamental constants since the beginning of time and the value of life today) that does not change the eigenvalue selected by the initial measurement (the value of the fundamental constants). As such, even if the search for the value of life by the universe’s evolution under the fundamental physical laws and constants were random, it would be endowed with quantum speedup and therefore teleological in character.
3.2 Formalization
We formalize the notion of the time-symmetric causal loops on the simplest, four drawer, instance of Grover quantum search algorithm. At the end of this subsection we will see how this immediately generalizes. The formalization consists of the following steps:
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem. We call this extended description the ordinary quantum description, of the quantum algorithm: ordinary because it is still a causal description that cannot describe causal loops.
(2) This point is propaedeutic to point (4). It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her full information about the solution of the problem444 We must assume that the problem-solver knows everything about the structure of the problem, thus also the relation between problem-setting and solution. .
(3) Complete the ordinary quantum description of the quantum algorithm by time-symmetrizing it.
(4) Relativize with respect to the problem-solver the time-symmetric causal loops generated by the time-symmetrization of point (3). This produces the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
3.2.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
The usual quantum description of quantum algorithms is limited to the process of solving the problem, where the quantum speedup resides, and lacks the initial measurement. Of course, it is trivially incomplete: the canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement. We extend the usual quantum description to the upstream process of setting the problem. An initial measurement of the problem-setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random. This is followed by the unitary evolution that computes the corresponding solution and the final measurement of the solution.
Let us introduce the notation. We make reference to the simplest four drawer instance of Grover algorithm:
We number the four draws in binary notation: $00,01,10,11$.
We will need two quantum registers of two quantum bits each:
A register $B$ under the control of the problem-setter Bob contains the problem-setting (the number of the drawer in which the ball is hidden). The content of register $B$ is the problem-setting observable $\hat{B}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{B}$ and $00$; $|01\rangle_{B}$ and $01$; $|10\rangle_{B}$ and $10$; $|11\rangle_{B}$ and $11$.
A register $A$ under the control of the problem-solver Alice contains the number of the drawer opened by her. The content of register $A$ is the observable $\hat{A}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$;
we call it content of $A$ observable. Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
The quantum description of the extended quantum algorithm, namely its ordinary quantum description is:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table I)
Here and in the following we disregard normalization.
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow555 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing. .
It is now Alice’s turn. The unitary part of her problem-solving action, $\hat{\mathbf{U}}_{1,2}$, produces the solution of the problem, of course, never changing the problem-setting. Note that, of $\hat{\mathbf{U}}_{1,2}$, it is sufficient to know that it does this, whereas we will never need to know how it does it. Of course, we must also know that there exists a unitary evolution that does this, it is the unitary part of Grover algorithm in the present case. Correspondingly, throughout this section, the unitary part of Grover algorithm will always remain just the symbol $\hat{\mathbf{U}}_{1,2}$.
At the end of Alice’s problem-solving action, at time $t_{2}$, register $A$ contains the solution of the problem, i.e. the number of the drawer with the ball – bottom-right corner of the diagram.
Eventually, by measuring $\hat{A}$ (the content of $A$), which is already in one of its eigenstates, Alice acquires the solution with probability $1$ – the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement. Correspondingly, there is a unitary transformation that connects the initial and the final measurement outcomes; the process between them is reversible despite comprising the final measurement (which does not change the quantum state).
3.2.2 Relativizing the ordinary quantum description of the quantum algorithm with respect to the problem-solver
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process. It is not a description for the problem-solver (Alice). In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement.
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by Bob and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following quantum description of the quantum algorithm with respect to Alice:
time $\begin{array}[]{ccc}\begin{array}[]{c}\text{time}\text{ }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{\text{time }\text{}}t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ &&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table II)
To Alice, the initial measurement of $\hat{B}$ leaves the initial state of register $B$ unaltered. At the beginning of her problem-solving action she is completely ignorant of the number of the drawer with the ball selected by Bob – see the state of register $B$ on the left of $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$.
The unitary transformation $\hat{\mathbf{U}}_{1,2}$ is then performed for a quantum superposition of the four possible numbers of the drawer with the ball. The state at the end of it – on the right of the horizontal arrows – is a quantum superposition of tensor products, each a possible number of the drawer with the ball in register $B$ tensor product the corresponding solution (that same number) in register $A$. Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution – bottom-right corner of Table II diagram.
Note that this relativization relies on relational quantum mechanics $\left[36,37\right]$, where the quantum state is relative to the observer. In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
3.2.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
As we have already said, in order to obtain the quantum superposition of the causal loops whose mutually exclusive $or$ is implied by the sheer existence of an optimal quantum speedup, we should first time-symmetrize the ordinary quantum description of the quantum algorithm, then relativize with respect to the problem-solver the causal loops obtained.
To start with, we should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes among all the possible pairs time-symmetrically shares between the initial and final measurements. The half information selected by the initial measurement should propagate forward in time, by $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement. Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time, by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$, until replacing the previous outcome of the initial measurement.
We consider the time-symmetrization instance where the initial measurement of $\hat{B}$ reduces to the measurement of $\hat{B}_{l}$, the left bit of the two-bit number contained in register $B$; say that it randomly selects left bit 00. The final measurement of $\hat{A}$ should correspondingly reduce to that of $\hat{A_{r}}$ the right bit of the two-bit number contained in register $A$; say that it randomly selects right bit $1$. The randomly selected number of the drawer with the ball is thus $01$.
In this particular time-symmetrization instance, the ordinary quantum description of the quantum algorithm with respect to the external observer of Table I becomes:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, }\text{}\text{meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }\text{}t_{2}\text{, meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}\\ &&\Downarrow\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table III)
The initial measurement of $\hat{B}_{l}$, selecting the 00 of $01$, projects the initial quantum superposition of all the basis vectors of register $B$ on the superposition of those beginning with 00 – vertical arrow on the left of the diagram. Under $\hat{\mathbf{U}}_{1,2}$, the latter superposition causally propagates forward in time into the superposition of the two tensor products on the right of the right looking horizontal arrows. Then the final measurement of $\hat{A_{r}}$, selecting the $1$ of $01$, projects the latter superposition on the term ending with $1$ under the right vertical arrow. Note that this outcome of the final measurement, which inherits both selections, coincides with that of the ordinary quantum description – see the bottom-right corner of Table I diagram. Time-symmetrically, this final measurement outcome causally propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation (which inherits both selections) is an instance of the time-symmetrized description of the quantum algorithm – bottom line of the diagram. Note that, since the symbol “$\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$” is mathematically equivalent to “$\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$”, this instance, up to the time-direction of causality, is the unitary process between the two one-to-one correlated measurement outcomes back again.
Also note that causality performs a causal loop that changes its own past; the outcome of its initial measurement of $\hat{B}_{l}$ is changed into the outcome of the initial measurement of $\hat{B}$ of the ordinary quantum description; compare the state of register $B$ under the left vertical arrow with the state in the bottom-left corner of Table III diagram.
We should eventually take the quantum superposition of all the time-symmetrization instances. We can see that, no matter how we evenly share the selection between initial and final measurements, the final measurement outcome, which inherits both selections, is always that of the final measurement of $\hat{A}$ of the ordinary quantum description of the quantum algorithm. Correspondingly, its subsequent backward in time propagation by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ (bottom line of the diagram), mathematically (disregarding the direction of causality), is always the ordinary quantum description of the unitary process between the two correlated measurement outcomes, namely the bottom line of Table I diagram.
Summing up, mathematically, time-symmetrization leaves the quantum description of the unitary process between the two measurement outcomes unaltered but changes the structure of causality along the process. The causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
3.2.4 Relativizing the causal loops with respect to the problem-solver
The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is relative to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup. This is because the latter causal loops, concerning the problem-solver advanced knowledge of one of the halves of the information about the solution, must of course be relative to the problem-solver.
Therefore, we should relativize with respect to the problem-solver the causal loop of Table III by postponing after her problem-solving action the projection of the quantum state due to the initial measurement of $\hat{B}_{l}$ (whose outcome should be concealed from Alice). This yields the causal loop in the following table:
time $\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }t_{2}\text{,meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ \begin{array}[]{c}\left(|01\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$ – top-left corner of the diagram – is postponed after Alice’s problem-solving action, namely outside the present table that is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in present assumptions selects the value $1$ of the right bit of the number contained in register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized description of the quantum algorithm with respect to the problem-solver.
Note that we would have obtained the identical diagram by time-symmetrizing the description of the quantum algorithm relativized to the problem-solver of Table II.
Also note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, the bottom line of the diagram of Table IV can also be read in the usual left to right way:
time $\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table V)
The entire causal loop is the zigzag diagram of table IV. It is a causal loop that changes its own past. The initial state in the top-left part of the diagram, where the state of register $B$ tells us that Alice is completely ignorant of the number of the drawer with the ball, changes into the initial state in the bottom-left part of the diagram, where the state of register $B$ tells us that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$. Since what happens along the time-symmetric causal loop obeys classical logic, the number of drawer openings needed to find the solution in the optimal quantum case is that needed in the optimal logical way, namely just one.
In other words, the computational complexity of the problem to be solved by her reduces. The problem of locating the ball in one of $4$ drawers – as for the state of register $B$ in the upper-left corner of the diagram of Table IV – becomes that of locating it in one of $2=\sqrt{4}$ drawers – as for the state of register $B$ in the bottom-left corner of Table IV diagram (or the left of Table V diagram). In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $4$ basis vectors to $2=\sqrt{4}$.
In equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces.
The generalization of all the above is straightforward. Of course, everything immediately generalizes to any number of drawers. Note that, in all our considerations, we never needed to know the unitary transformation ($\hat{\mathbf{U}}_{1,2}$) between the two measurement outcomes. It sufficed to know that there exists a unitary transformation that changes its input state, where the two observables $\hat{B}$ and $\hat{A}$ are unrelated, into its output state, where they are one-to-one correlated, without changing the eigenvalue of the observable selected by the initial measurement.
4 The possible physical basis of natural teleological evolutions
We have seen in the Introduction how, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward the value of life (in particular, either sentient life or absence of life) under the fundamental physical laws and constants has the form of a quantum search algorithm. As such, it should be endowed with quantum speedup and would therefore be teleological in character even in the case that the search for the value of life is assumed to be random.
In the universe we are in, the search for the value of life becomes search for life, of course endowed with the same quantum speedup. We have also seen that, as a consequence, the evolution of life on Earth, obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speed up of the evolution of the universe toward life and would therefore be teleological too. The same would necessarily hold for the evolutions of species, etc..
In this section, after encoding a toy model of the evolution of the universe toward the value of life on Grover quantum search algorithm, we will discuss the many impacts that the present physical notion of teleological natural evolution should have on different areas of science and knowledge.
4.1 Teleological evolution of the universe toward the value of life – a toy model
A toy model of a teleological evolution of the universe toward the value of life can be trivially mapped on the simplest instance of Grover algorithm.
We should imagine a value of the fundamental physical constants observable $\hat{B}$ and a value of life observable $\hat{A}$. Say that the eigenvalue $01$ of $\hat{B}$ is the value of the fundamental physical constants compatible with the development of sentient life and the eigenvalues $00,10,11$ the values of those incompatible. We assume that, correspondingly, the eigenvalue $01$ of the measurement of $\hat{A}$ at the present time would mean existence of sentient life and the eigenvalues $00,10,11$ absence of life.
Initially, $\hat{B}$ should be in a quantum superposition (or, indifferently, mixture) of all its eigenstates (i.e of all the possible universes each with its fundamental physical constants). We can assume that also $\hat{A}$, the value of life observable, is initially in an independent quantum superposition of all its eigenstates.
The universe’s unitary evolution under the fundamental physical laws and constants would have finally, at the present time, built up a quantum superposition of tensor products, each an eigenstate of $\hat{B}$ tensor product the corresponding eigenstate of $\hat{A}$ – see Table II. In our universe, at the present time one with observers, the measurement of $\hat{A}$ at the present time would select the eigenvalue $01$ of $\hat{A}$, namely “existence of sentient life” and the corresponding eigenvalue $01$ of the value of the fundamental physical constants observable $\hat{B}$. Since the universe’s evolution does not change the value of the fundamental physical constants, all is as if the eigenvalue of $\hat{B}$ was selected by a measurement of $\hat{B}$ at the beginning of time, when there could be no observers.
Of course, this is exactly the quantum search algorithm described in the previous section, the version with respect to the problem-solver of Table II. As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup, and does even if the quantum search for the value of sentient life is random.
We should note that the present toy model is less crude than it may seem. In fact, it immediately generalizes to any number $N$ of drawers. This means that any amount of information can be encoded in the process we are dealing with. Moreover, the unitary transformation $\hat{\mathbf{U}}_{1,2}$ between the initial and final one-to-one correlated measurement outcomes is any unitary transformation that does not change the eigenvalue selected by the initial measurement. Therefore, in principle, this toy model could even become a faithful discrete model of the actual evolution of the universe toward the value of life.
4.2 Decohered teleological evolutions: what would they look like?
The supposedly teleological evolutions of life would be, so to speak, around us in the macroscopic world. Decohered and endowed with quantum speedup it is natural to wonder what they would look like. Now, all natural evolutions are necessarily described in a probabilistic way. We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical stochastic process, transition probabilities can depend on the past of the process but not on the future of it.
Instead, a decohered process teleological in character must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection because more fit to the environment must make, when they occur in the present, the transition toward them more probable. By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the evolution would be difficult to distinguish from a classical stochastic evolution. Let us call time-symmetric the present teleological kind of stochastic evolution, given that its description would be generated by a time-symmetrization procedure.
Let us mention what could be an emblematic example of time-symmetric stochastic evolution. It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
The usual explanation is of course a causal one: at first, feathers appeared by random mutation and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
According to the present teleological view of the evolutions of the living, even during the flightless period of the feathered dinosaurs, those random DNA mutations that would have led in the future to the conquest of the air – obviously an extremely successful outcome from the point of view of survival – would have more likely occurred.
The above explanation is puzzling, of course. However, let us note that it is not more puzzling than the complete quantum description of the quantum computational speedup. We should keep in our mind that we are under the quantum cosmological assumption that should make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as puzzling as the microscopic quantum world.
4.3 How to model teleological natural evolutions
We address the problem of how to model a decohered evolution endowed with quantum speedup – say that of the living on Earth after the beginning of the liquid water stage of it. The canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest. Although, in the present case, this obviously holds only in principle, it would be interesting to apply it to a toy model of the situation.
One should identify the purely quantum evolution of a system (standing for the universe) endowed with quantum speedup, where the achievement of the goal is mostly localized in a subsystem thereof (standing for the Earth). Then, one should obtain a decohered evolution of the subsystem from that of the system by tracing out all the rest. Note that the interaction of the subsystem with the rest of the system, in the present case, would do the opposite than decoherence. It would be what allows the evolution of the subsystem toward the goal to occur with quantum speedup as part of the purely quantum evolution of the system.
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[38\right]$.
As already said, a further possibility should be to model a decohered evolution endowed with quantum speedup as a stochastic process with transition probabilities time-symmetrically depending on both the past and the future of the process, something just outlined for the time being.
4.4 Impact on the natural sciences
The current position of the natural sciences toward teleology can be considered in some measure conflicting. On the one hand, the teleological view would seem to allow a deeper understanding of the functional structure and the evolutions of living beings. On the other, as this vision is considered without physical basis and therefore misleading, there is outright hostility to it.
This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[39\right]$:
Teleology is like a mistress to a biologist: he cannot live without her, but he is unwilling to be seen with her in public.
It would seem that, today, no evolutionary biologist supports the teleological view, some explicitly assert that it is an unworkable view because it implies retrocausality.
No biologist, of course, does not mean no scientist. A paradigmatic example is the American philosopher Nagel $\left[40\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
We would put together Nagel’s view in support of the teleological character of the evolutions of the living and Haldane’s assertion that a biologist cannot live without teleology to affirm that the notion of teleological natural evolution is more “scientific” than the mainstream one of causal natural evolution. This in the sense that it better answers empirical evidence – of course the first canon of science. The only reason biologists would be against teleology despite empirical evidence would come from another science. We mean the notion – current in today’s physics – that a teleological evolution is physically impossible.
Of course, the sheer notion that quantum computation is teleological in character may already put a strain on the position of today’s biology: actually, although in the quantum world, teleological evolutions are physically possible. The strain might increase with the notion that, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward the value of life, toward life in the universe we are in, should be teleological in character. It should further increase with the consequent notion that also the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, even if decohered, should inherit the teleological character of the universe’s evolution.
Obviously, all this would go against the ban currently placed by the natural sciences on the teleological view.
Instead, it might go along with the fact that, in recent years, a group of evolutionary biologists $\left[41\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a possible explanation.
It would also go along with the fact that some scientists have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[38\right]$.
4.5 Impact on philosophy
Of course, the current theory of teleological natural evolutions has a precise conceptual precedent in a very ancient notion of natural philosophy $\left[42\right]$. We should therefore be allowed to look for other possible cases where philosophical thinking has anticipated aspects of it.
In effect, the teleological theory in question seems to revamp a full body of ancient philosophical views sort of opposite (and complementary) to the body of philosophical views emerged with the scientific revolution.
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
The causal view and the mutual causality view seem to have respectively subtended two opposing school of thought throughout the history of philosophy. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views respectively.
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[32\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$ is a mutual causality view.
We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond, respectively, to the causal and mutual causality views of evolutions.
Regarding reductionism vs. monism, we note that deriving the teleological character of the decohered evolutions of living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view. It presupposes that the description of what we see around and within us is essentially inseparable from the description of the whole.
Let us mention other theories in the history of philosophy reminiscent of the mutual causality between sentient life today and the physical laws with their physical constants since the beginning of time:
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[43\right]$. In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. The identity between the fundamental physical laws since the beginning of time and the ideas of our mind today is formally reminiscent of the identity (and mutual causality) between the number of the drawer with the ball (the problem-setting) randomly selected by Bob’s measurement at the beginning of the quantum algorithm and the number of the drawer with the ball (the solution of the problem) selected by Alice’s measurement at the end of it. It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, the fact it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
The above should also go along with Fritjov Capra’s $\left[44\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
(ii) We do not need to be anthropocentric. This mutual causality would also be between fundamental physical laws and constants on the one side and spatial intelligence, often superior to our in non-human animals, on the other. The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality). Indeed a quantum speedup in the development of spatial intelligence – and the consequent teleological character of this development – would imply the mutual causality in question.
(iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[45\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
(iv) It is natural to wander whether the present time-symmetric causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
(v) Michelangelo’s “Creation of Adam” on the ceiling of the Cappella Sistina could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[46\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today?
To sum up, it seems clear that the contrast between the causal and mutual causality views corresponds to an opposition that characterizes the entire history of thought. We would ascribe this unified view of the history of philosophy to the unifying power of the present “scientific” notion of teleological evolution that, under quantum cosmology, also extends to the macroscopic world.
5 Research prospects
Future work on the subject of the present work could have a broad spectrum.
For what concerns quantum physics:
(i) Look for a synthesis between the present notion of the time-symmetric causal loops and time-symmetric quantum mechanics $\left[19-23\right]$ to which it is obviously inspired.
(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem, although decohered, inherits the quantum speedup and the consequent teleological character of the evolution of a purely quantum system. Of course, here, the quantum system stands for the universe and the subsystem for the Earth subsystem.
Note that the present notion for which the decohered behavior of a subsystem inherits quantum features from the behavior of the quantum system might apply to other cases. Possible examples are the life processes suspected to have higher than classical efficiency and consciousness suspected to be endowed with quantum nonlocality, the quantum system being here, of course, a quantum universe.
(iii) Look for a possible impact of the present teleological view on quantum cosmology. According to current assumptions, even the “inanimate” part of the universe’s evolution, an evolution toward life too, should be endowed with quantum speedup and thus have teleological character.
(iv) The Anthropic Principle is limited to asserting the correlation between the “value” of life today and the value of the fundamental physical constants since the beginning of time. One could also look for a relation between the value of life today and the character of the fundamental physical laws since the beginning of time. Of course, in the universe we are in, also the latter character should have an attractor in the goal of creating life.
About the natural sciences, looking for:
(I) A reconsideration of the ban on applying the teleological view and language to natural evolutions. Restoring the dismissed notion of teleological evolution would obviously be a paradigm shift for today’s natural sciences.
(II) Toy models of teleological Darwinian evolutions. These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
Note that there might be a divide between the teleological view of evolution and the, of course dominant, causal/classical view. An evolution endowed with quantum speedup might not only mean a faster evolution; the mutual causality implicit in it could produce forms of evolution that are precluded to the causal/classical type.
(III) An application of the time-symmetric causal loops to quantum self-organization. An example could be studying the teleological formation of a self-replicating system. There might be an enormous jump of complexity in the transition from a soup of life-precursor molecules to a self-replicating system. The idea that this transition has an attractor in the full fledged self-replicating system can therefore be attractive.
Finally, opposing schools of thought throughout the history of philosophy could be reviewed in the light of the opposition between the causal and mutual causality views, as exemplified in the previous subsection.
6 Conclusions
It might be useful to summarize the key points of our unconventional argument.
We have (trivially) shown that the sheer existence of the quantum speedup of the optimal Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
The following is the consequent line of thinking.
Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic of the logical implication in question. Therefore, Grover algorithm must host the quantum superposition of the causal loops we are dealing with – logically mutual exclusive possibilities go in quantum superpositions in the quantum world. However, of course, this cannot be in the quantum description of the quantum algorithm that, being causal, cannot describe causal loops.
It should be noted that we adopt a canonical quantum description of quantum processes, consisting of initial measurement, unitary evolution, and final measurement. In quantum algorithms, the initial measurement selects the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings, the unitary evolution computes the corresponding solution, and the final measurement reads it.
Time-symmetrizing this, still causal, ordinary quantum description of the quantum algorithm generates the quantum superposition of the causal loops logically implied by the sheer existence of its quantum speedup.
We require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes; that the selection performed by the initial measurement propagates forward in time until becoming the state immediately before the final measurement and the selection performed by the final measurement propagates backward in time until replacing the previous outcome of the initial measurement. This is a causal loops that changes its own past – the initial measurement outcome. It is indeed one of the causal loops logically implied by the sheer existence of the quantum speedup.
Of course, we should take the quantum superposition of all the possible time-symmetrization instances, each for each possible way of time-symmetrically sharing the selection of the pair of measurement outcomes between the initial and final measurements. This leaves the quantum description of the unitary evolution between the initial and final measurement outcomes mathematically unaltered but changes the behavior of causality along the evolution. The single causal process of the ordinary quantum description is replaced by the quantum superposition of causal loops logically implied by the sheer existence of the quantum speedup. Being generated by a time-symmetrization, we call these causal loops time-symmetric.
Note that the time-symmetrization in question turns out to be the only possible way of obtaining a quantum description of the quantum algorithm with the quantum superposition of causal loops it must host. It is therefore mandatory.
We have shown that this immediately generalizes to any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (the problem-setting in quantum algorithms). The ordinary quantum description of any such process is incomplete and is completed by time-symmetrizing it. Let us repeat ourselves: this leaves the ordinary quantum description of the unitary part of the quantum process mathematically unaltered but changes the description of the behavior of causality along that part. The single causal process of the ordinary quantum description is replaced by a quantum superposition of time-symmetric causal loops.
In particular, the generalization covers the case of quantum nonlocality, where the unitary evolution is the process of spatial separation of the two subsystems in a state of maximal entanglement between the respective observables. There is one-to-one correlation between the outcomes of measuring these observables in the same basis. Here, the time-symmetric quantum causal loops play the role of the hidden variables envisaged by Einstein and the others to complete the ordinary quantum description of quantum nonlocality.
Eventually, the quantum causal loop description of quantum algorithms highlights their teleological character, that is, their being evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again). This naturally raises the question whether the time-symmetric causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
It can be argued that investigating this possibility is incumbent on the part of the physical science toward the natural sciences, after its having asserted for centuries that teleological evolutions are unphysical.
This immediately raises the problem of explaining the presence of quantum features (the quantum speedup) in evolutions of the living that can be the size of the Earth and have durations of billions of years. The obvious impossibility that the quantum features of the microscopic world scale up to such levels suggests the only possible solution: that they scale down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time, we should assume, with quantum cosmology, that the evolution of the universe is unitary quantum.
Surprisingly, the “Fine-tuned Universe” version of the Anthropic Principle puts all the above things together. It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”. We assume for simplicity that this correlation is one-to-one.
The assumption that the evolution of the universe is unitary quantum, of course forces us to adopt a quantum interpretation of the Anthropic Principle. It is now the unitary evolution of the universe that builds up a quantum correlation between the value of the physical constants since the beginning of time and the value of life today.
We have seen that any unitary evolution that builds up one-to-one correlation between an initial and a final measurement outcome without changing the eigenvalue selected by the initial measurement can occur with quantum speedup and does even if the search for that correlation is random. That the search for the correlation is random is of course a conservative assumption; that there has been this search by the evolution of the universe is what the Anthropic Principle says.
In conclusion, the unitary evolution of the universe toward the value of life occurs with quantum speedup and is thus teleological. In the universe we are in, the same, of course, holds for the evolution of the universe toward life.
The evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out all the rest, should inherit the teleological character of the latter evolution. In fact: (i) evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe and (ii) the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
In conclusion, we have seen that the ordinary, causal, quantum description of any unitary evolution between an initial and a final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it. This generates a quantum superposition of time-symmetric causal loops that explain such diverse quantum phenomena as the quantum speedup and the consequent teleological character of quantum algorithms, quantum nonlocality, and, under a quantum cosmological interpretation of the Anthropic Principle, the teleological character of the evolutions of life.
Although the time-symmetric causal loops we have been dealing with should be exempt from the paradoxes that plague the usual notion of causal loop insofar as they are imposed by logic and quantum mechanics, they remain puzzling for the consequences they entail. Being related to quantum correlation, they can also be thought to be pervasive. Once verified, their existence could introduce a paradigm shift in quantum mechanics.
Acknowledgments
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments throughout the whole of the present retrocausal approach to the quantum speedup, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the Elsag Bailey-ISI Turin Villa Gualino workshops of the years 1991-96.